OpenStudy (anonymous):

partial derivative with respect to y- fy(x,y)= sqrt 4x^2 - 5xy^3 + 12y^2x^2??

OpenStudy (badhi):

$$f(x,y)=\sqrt{4x^2-5xy^3-12y^2x^2}$$ consider x as a constant and do the normal differentiation wrt y

OpenStudy (anonymous):

i haven't done this kind of math in a long time... can u show me the steps??

OpenStudy (badhi):

Are you familiar with single variable differentiation?

OpenStudy (anonymous):

uh sorta... it's f(x+h) - f(x)/ h right?

OpenStudy (badhi):

that's the basis, I meant the proven things that we normally apply directly for differentiation. such as $$\frac{d x^n}{dx}=nx^{n-1}$$ and some other rules such as chain rule How about them?

OpenStudy (badhi):

You should be quite comfortable with single variable calculus to handle multi-variate differentiation. Other wise I recommend that you should first remind the things in single variable calculus

OpenStudy (anonymous):

ok thank you i'll go from there!

OpenStudy (badhi):

Since you've asked, I'll show the steps for your question since the differentiation is respect to y we can consider x as a constant thus the expression is like, $$f(y)=\sqrt{a + by^3 + cy^2}$$ where a,b,c are considered constants (a=4x^2, b=-5x, c=-12x^2) if we take $$v=a+by^3+cy^2$$ $$f(v)=\sqrt{v}$$ Now we have to do the differentiation. For that we use chain rule i.e. $$\frac{du}{dy}=\frac{du}{dv}\frac{dv}{dy}$$ here we substitute u with f, $$\frac{df}{dy}=\frac{df}{dv}\frac{dv}{dy}$$ This means that we need to find df/dv and dv/dy to find df/dy $$\frac{df}{dv}=\frac{d\sqrt{v}}{dv}=\frac{1}{2\sqrt{v}}$$ $$\frac{dv}{dy}=\frac{d(a+by^3+cy^2)}{dy}=3by^2+2cy$$ so, $$\frac{df}{dy}=\frac{1}{2\sqrt{v}}(3by^2+2cy)=\frac{3by^2+2cy}{2(a+by^3+cy^2)}=\frac{3(-5x)y^2+2(-12x^2)y}{2(4x^2-5xy^3-12y^2x^2)}$$

OpenStudy (badhi):

the last line should be changed to, $$\frac{df}{dy}=\frac{3by^2+2cy}{2\sqrt{a+by^3+cy^2}}=\frac{3(-5x)y^2+2(-13x^2)y}{2\sqrt{4x^2-5xy^3-12x^2y^2}}$$ I forgot the square root :(